OpenStudy (anonymous):

HELP i need help with my algebra can some one please help me i will medal fan and follow on ig

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

@rosecute

OpenStudy (anonymous):

@robtobey

OpenStudy (midhun.madhu1987):

question?

OpenStudy (anonymous):

\[11+4b=8b-13\]

OpenStudy (midhun.madhu1987):

4b - 8b = -13 - 11

OpenStudy (midhun.madhu1987):

combine thema nd find b

OpenStudy (anonymous):

i dont now how

OpenStudy (midhun.madhu1987):

4-8 = ?

OpenStudy (anonymous):

-4

OpenStudy (midhun.madhu1987):

so 4b-8b=?

OpenStudy (anonymous):

-4b

OpenStudy (midhun.madhu1987):

yes... and -13-11=?

OpenStudy (anonymous):

-24

OpenStudy (midhun.madhu1987):

okay... so -4b = -24 now divide both sies by -4 and u get b

OpenStudy (anonymous):

what is it

OpenStudy (midhun.madhu1987):

-4b = -24 divide both sides of the equation by -4.

OpenStudy (anonymous):

one one side i get 7 and another i get 1 now what

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (midhun.madhu1987):